### v1

# Generalizing Goodstein’s theorem and Cichon’s independence proof

Abstract
We generalize Goodstein’s theorem [2] and Cichon’s independence proof [1] to $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$ using results from [3]. The method is generalizable to stronger notation systems that provide unique terms for ordinals and enjoy Bachmann property.
1 Introduction
Let $\mathcal{E}$} be the familiar notation system for ordinals below $\varepsilon_{0}$, the proof-theoretic ordinal of Peano arithmetic $\operatorname{PA}$}, built from 0, $(\xi,\eta)\mapsto\xi+\eta$, and $\xi\mapsto\omega^{\xi}$ using Cantor normal form, where $\omega$ denotes the least infinite ordinal number. For $\alpha\in{\mathcal{E}}$ let ${\operatorname{mc}}(\alpha)$ be the maximum counter of multiples that occur in $\alpha$, that is, the maximum number of times the same summand is consecutively added in the notation of $\alpha$. For $k\in[2,\omega)$ define the quotient $\nicefrac{{{\mathcal{E}}}}{{k}}$$\subseteq{\mathcal{E}}$ by
[…]
Assuming the familiar system $({\mathcal{E}},\cdot[\cdot])$ of fundamental sequences (limit approximations) for $\mathcal{E}$} and the definition of the slow-growing hierarchy $\mathrm{G}_{k}$} from [1], given by
${\mathrm{G}_{k}}(0):=0,\>{\mathrm{G}_{k}}(\alpha+1):={\mathrm{G}_{k}}(\alpha)+1,\mbox{ and }{\mathrm{G}_{k}}(\lambda):={\mathrm{G}_{k}}(\lambda[k])\mbox{ for }\lambda\in{\mathcal{E}}\cap\mathrm{Lim},$ (1)
as shown in [1] the restricted mapping
[…]
is a bijection via transforming back and forth the bases $k$ and $\omega$, that is, writing a natural number $N$ in hereditary base-$k$ representation using 0, addition, and exponentiation to base $k$, replacing the base $k$ by $\omega$ we obtain ${\mathrm{G}_{k}^{-1}}(N)$, the unique preimage in $\nicefrac{{{\mathcal{E}}}}{{k}}$}. The image of $\nicefrac{{{\mathcal{E}}}}{{k}}$} under $\mathrm{G}_{k}$} is the Mostowski collapse of $\nicefrac{{{\mathcal{E}}}}{{k}}$} and equal to $\omega$, i.e., the restriction of $\mathrm{G}_{k}$} to $\nicefrac{{{\mathcal{E}}}}{{k}}$} is an order isomorphism.
Since $({\nicefrac{{{\mathcal{E}}}}{{k}}})_{k\in[2,\omega)}$ is $\subseteq$-increasing, base transformation from $k$ to $l$ for $2\leq k<l<\omega$ is characterized by
$N[k\mapsto l]={\mathrm{G}_{l}}(N[k\mapsto\omega]),$
where $N<\omega$ is assumed to be given in hereditary base-$k$ representation, so that $N[k\mapsto\omega]\in{\nicefrac{{{\mathcal{E}}}}{{k}}}\subseteq{\nicefrac{{{\mathcal{E}}}}{{l}}}$. Thus, ordinals $\alpha\in{\mathcal{E}}$ can be seen as direct limit representations of natural numbers in hereditary base-$k$ representation, provided that ${\operatorname{mc}}(\alpha)<k$.
In this article we generalize this mechanism to the notation system $\operatorname{T}$} for ordinals below Takeuti ordinal, that is, to strength $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$, using fundamental sequences and machinery introduced in [3]. The method is generalizable to stronger notation systems that provide unique terms for ordinals and enjoy Bachmann property.
Adopting the preliminaries of [3], subsections 2.1, 2.3, and 2.5, which are summarized here in subsection 2.1 for reader’s convenience, let $\operatorname{T}$} be the notation system $\operatorname{T}^{\tau}$} for $\tau=1$ introduced there and let $\mathring{{\operatorname{T}}}$} be the subset of terms of countable cofinality. We called the system $({\mathring{{\operatorname{T}}}},\cdot[\cdot])$ of fundamental sequences with Bachmann property given in Definition 3.5 of [3] a Buchholz system. Bachmann property is a nesting property of fundamental sequences, namely that for terms $\alpha$ and $\beta$ such that $\alpha[\zeta]<\beta<\alpha$ and that are not regular cardinals, we have $\alpha[\zeta]\leq\beta[0]$, cf. Lemma 4.1 and Theorem 4.2 of [3].
For compatibility reasons with [1] we need to modify the definition of fundamental sequences in [3] so that for ordinals in the class of additive principal numbers, i.e. ordinals in the image of $\omega$-exponentiation $\xi\mapsto\omega^{\xi}$, say $\eta$, we redefine $(\eta\cdot\omega)[k]:=\eta\cdot k$, rather than $(\eta\cdot\omega)[k]:=\eta\cdot(k+1)$. [^1] Note that with this modification Bachmann property then only holds with appropriate restriction, which we will address later.
We will be able to use the approach by Cichon [1], transferring the definition of collapsing functions $\mathrm{G}_{k}$} as defined in (1) and predecessor functions $\operatorname{P}_{k}$}, defined as in [1] by
${\operatorname{P}_{k}}(0):=0,\>{\operatorname{P}_{k}}(\alpha+1):=\alpha,\mbox{ and }{\operatorname{P}_{k}}(\lambda):={\operatorname{P}_{k}}(\lambda[k])\mbox{ for }\lambda\in\mathrm{Lim},$ (2)
to the (modified) Buchholz system $({\mathring{{\operatorname{T}}}},\cdot[\cdot])$, more specifically to its countable initial segment. Note first of all that Cichon’s crucial lemma also holds in our context, with the same proof by induction on $\alpha$:
Lemma 1.1 (cf. Lemma 2 of [1])
For $k\in[2,\omega)$ and $\alpha\in{\operatorname{T}}\cap\Omega_{1}$ the operations $\mathrm{G}_{k}$} and $\operatorname{P}_{k}$} commute, thus
${\mathrm{G}_{k}}{\operatorname{P}_{k}}\alpha={\operatorname{P}_{k}}{\mathrm{G}_{k}}\alpha,$
where the latter is equal to ${\mathrm{G}_{k}}\alpha-1$.
We are going to introduce quotients $\nicefrac{{{\operatorname{T}}}}{{k}}$} and $\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}$} of $\operatorname{T}$} and $\mathring{{\operatorname{T}}}$}, respectively, to base $k$ where $k\in[2,\omega)$. The sets ${{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}:={\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}}\cap\Omega_{1}$, each of which is cofinal in ${\operatorname{T}}\cap\Omega_{1}$ and which are $\subseteq$-increasing in $k$, will be the canonical analogues of the quotients $\nicefrac{{{\mathcal{E}}}}{{k}}$} mentioned above. Consider as an instructive basic example the tower $k_{k}$ of height $k$ of exponentiation to base $k$. Clearly, $k_{k}[k\mapsto\omega]\in{\nicefrac{{{\mathcal{E}}}}{{k}}}$, so that the direct limit corresponding to $k_{k}$ in $\mathcal{E}$} is $\omega_{k}$, the tower of height $k$ of exponentiation to base $\omega$. This does not hold in $\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}$}, where the corresponding ideal object is $\varepsilon_{0}=\lim_{n<\omega}\omega_{n}$. The sets ${\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}$} therefore provide more refined hierarchies of ideal objects uniquely denoting natural numbers, and functions $\mathrm{G}_{k}$} will act as their enumeration functions, see Theorem 2.14.
These preparations allow us to quite canonically generalize Goodstein’s theorem [2] and Cichon’s independence proof in [1] to obtain independence of the theory $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$, the biggest of the “big five” theories in Simpson’s book on reverse mathematics. The argumentation, “Cichon’s trick”, is as follows: Given a starting base $k\in[2,\omega)$ and a natural number $N\in{\mathbb{N}}$, by Theorem 2.14 we obtain a unique $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$ such that
$N={\mathrm{G}_{k}}(\alpha)=:N_{1}.$
The Goodstein operation of incrementing the base of representation and subsequent subtraction of $1$ is then given by
$N_{1}[k\mapsto k+1]-1={\mathrm{G}_{k+1}}(\alpha)-1={\operatorname{P}_{k+1}}{\mathrm{G}_{k+1}}\alpha={\mathrm{G}_{k+1}}{\operatorname{P}_{k+1}}\alpha=:N_{2},$
where we used Lemma 1.1. Iterating the procedure yields
$N_{2}[k+1\mapsto k+2]-1={\mathrm{G}_{k+2}}({\operatorname{P}_{k+1}}\alpha)-1={\operatorname{P}_{k+2}}{\mathrm{G}_{k+2}}({\operatorname{P}_{k+1}}\alpha)={\mathrm{G}_{k+2}}{\operatorname{P}_{k+2}}({\operatorname{P}_{k+1}}\alpha)=:N_{3},$
so that we obtain the generalized Goodstein sequence
$N_{l+1}={\mathrm{G}_{k+l}}({\operatorname{P}_{k+l}}\ldots{\operatorname{P}_{k+1}}\alpha).$
Termination of the generalized Goodstein sequence is therefore expressed by
$\exists l\>N_{l+1}=0,$
and noting that ${\mathrm{G}_{k}}\alpha=0$ if and only if $\alpha=0$, we may reformulate the generalized Goodstein principle as follows:
$\forall k\in[2,\omega)\>\forall\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}\>\exists l\>\>{\operatorname{P}_{k+l}}\ldots{\operatorname{P}_{k+1}}\alpha=0.$ (3)
Thus, transfinite induction up to Takeuti ordinal proves this generalized Goodstein principle. Independence then follows, once we show that the function ${\mathrm{h}_{k}}:{\operatorname{T}}\cap\Omega_{1}\to{\mathbb{N}}$ defined by
${\mathrm{h}_{k}}(\alpha):=k+\min\{l\mid{\operatorname{P}_{k+l}}\ldots{\operatorname{P}_{k+1}}\alpha=0\}$ (4)
can easily be expressed in terms of the Hardy function $H_{\alpha}$, as in [1] for the original Goodstein principle. Defining Hardy hierarchy along Takeuti ordinal (i.e. the initial segment ${\operatorname{T}}\cap\Omega_{1}$) in the same way as in [1] by
$H_{0}(x):=x,\>H_{\alpha+1}(x):=H_{\alpha}(x+1),\mbox{ and }H_{\lambda}(x):=H_{\lambda[x]}(x),$ (5)
we obtain by straightforward <<-induction on $\alpha$:
${\mathrm{h}_{k}}(\alpha)=H_{\alpha}(k).$ (6)
2 Ordinal algebra
2.1 Preliminaries
We denote the class of limit ordinals as $\mathrm{Lim}$, the class of additive principal numbers as $\mathbb{P}$}, and the class of (non-zero) ordinals closed under $\omega$-exponentiation, also called $\varepsilon$-numbers, by $\mathbb{E}$}.
For terms $\alpha$ we use two abbreviations: first, writing $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$ means that $\eta\in{\mathbb{P}}$ and $\xi>0$ is minimal such that $\alpha=\xi+\eta$, second, $\alpha=_{\mathrm{\scriptscriptstyle{ANF}}}\xi_{1}+\ldots+\xi_{k}$ means that $\xi_{1},\ldots,\xi_{k}\in{\mathbb{P}}$ and $\xi_{1}\geq\ldots\geq\xi_{k}$. To indicate Cantor normal form representation of an ordinal $\alpha$, we write $\alpha=_{\mathrm{\scriptscriptstyle{CNF}}}\omega^{\xi_{1}}+\ldots+\omega^{\xi_{k}}$ if $\xi_{1}\geq\ldots\geq\xi_{k}$ and $k<\omega$. For the result of $l$-fold addition of the same term $\eta$ to a term $\xi$, where $l<\omega$, we will sometimes use the shorthand $\xi+\eta\cdot l$.
Terms in $\operatorname{T}$} are built up from 0, ordinal addition, and functions $\vartheta_{i}$ where $i<\omega$, where $\vartheta_{i}(0)={\Omega_{i}}$ and arguments of $\vartheta_{i}$ are restricted to terms below $\Omega_{i+2}$, setting $\Omega_{0}:=1$ and $\Omega_{i+1}:=\aleph_{i+1}$. As in subsection 2.1 of [3], slightly abusing notation we may consider notation systems ${\operatorname{T}}^{\Omega_{i+1}}$ to be systems relativized to the initial segment $\Omega_{i+1}$ of ordinals and built up over $\Omega_{i+1}=\vartheta_{i+1}(0),\Omega_{i+2}=\vartheta_{i+2}(0),\ldots$, i.e. without renaming the indices of $\vartheta$-functions. In this terminology ${\operatorname{T}}={\operatorname{T}}^{\Omega_{0}}$, and the sequence $({\operatorname{T}}^{{\Omega_{i}}})_{i<\omega}$ can be seen as an increasing sequence with larger and larger initial segments of ordinals serving as parameters.
The operation $\cdot^{\star_{i}}$ searches a $\operatorname{T}$}-term for its $\vartheta_{i}$-subterm of largest ordinal value, but under the restriction to treat $\vartheta_{j}$-subterms for $j<i$ as atomic. If such largest $\vartheta_{i}$-subterm does not exist, $\cdot^{\star_{i}}$ returns 0. The functions $\vartheta_{i}$ are natural extensions of the fixed point free Veblen functions and hence injective with images contained in the intervals ${\Omega_{i}},\Omega_{i+1})$, respectively for $i<\omega$. For $\alpha=\vartheta_{i}(\xi)$ and $\beta=\vartheta_{i}(\eta)$ (where $\xi$ and $\eta$ are any elements of the domain of $\vartheta_{i}$) we have
$\alpha<\beta\Longleftrightarrow(\xi<\eta\mbox{ and }\xi^{\star_{i}}<\beta)\mbox{ or }\alpha\leq\eta^{\star_{i}}.$
As in [3] the arguments of $\vartheta_{i}$-terms $\vartheta_{i}(\xi)$ are often written in a form $\vartheta_{i}(\Delta+\eta)$, i.e. $\xi=\Delta+\eta$. This by convention always indicates that $\Omega_{i+1}\mid\Delta$ (possibly $\Delta=0$) and $\eta<\Omega_{i+1}$. Informally, we sometimes call the multiple $\Delta$ of $\Omega_{i+1}$ the fixed point level of $\vartheta_{i}(\xi)$, and the predicate $F_{i}(\Delta,\eta)$ means that $\eta=\sup_{\sigma<\eta}\vartheta_{i}(\Delta+\sigma)$, which according to Proposition 2.6 of [3] holds if and only if $\eta$ is of a form $\vartheta_{i}(\Gamma+\rho)$ where $\Gamma>\Delta$ and $\eta>\Delta^{\star_{i}}$.
Any $\xi\in{\operatorname{T}}\cap{\mathbb{P}}$ is of a form $\vartheta_{i}(\Delta+\eta)$ for unique $i<\omega$, $\Delta\in{\operatorname{T}}\cap\Omega_{i+2}$ with $\Omega_{i+1}\mid\Delta$, and $\eta<\Omega_{i+1}$. We will sometimes write $\xi\cdot\omega$ as a shorthand for the additive principal successor of $\xi$ in $\operatorname{T}$}, instead of $\vartheta_{i}(\xi)$ in case of $\Delta>0$ and $\vartheta_{i}(\eta+1)$ if $\Delta=0$.
For terms $\alpha$ of a form $\vartheta_{i}(\Delta+\eta)$ the $\Omega_{i}$}-localization, cf. subsection 2.3 of [3], is a sequence ${\Omega_{i}}=\alpha_{0},\ldots,\alpha_{m}=\alpha$ ($m\geq 0$), in which for $\alpha>{\Omega_{i}}$ the terms $\alpha_{i}=\vartheta_{i}(\Delta_{i}+\eta_{i})$, $i=1,\ldots,m$, provide a strictly increasing sequence of $\vartheta_{i}$-subterms of $\alpha$ that are not in the scope of any $\vartheta_{j}$-function with $j<i$ (i.e. that are not collapsed), such that the sequence of fixed point levels $\Delta_{1},\ldots,\Delta_{m}$ is strictly decreasing and of maximal length. Localization therefore approximates the ordinal $\alpha$ from below in terms of decreasing fixed point levels.
Restricting ordinal addition occuring in $\operatorname{T}$} to summation in additive normal form, the terms in $\operatorname{T}$} uniquely identify ordinal numbers below Takeuti ordinal. The term algebra developed in [3] and here therefore is algebra of ordinals.
For reader’s convenience we recall the following simple but crucial definition of a characteristic function regarding uncountable moduli and then provide the concrete system of fundamental systems which is a slightly modified version of Definition 3.5 of [3] that fits our context.
Definition 2.1 (3.1 and 3.3 of [3])
We define a characteristic function $\chi^{\Omega_{i+1}}:{\operatorname{T}}^{\Omega_{i+1}}\to\{0,1\}$ where $i<\omega$ by recursion on the build-up of ${\operatorname{T}}^{\Omega_{i+1}}$:
  1. 1.
    $\chi^{\Omega_{i+1}}(\alpha):=\left\{\begin{array}[]{cl}0&\mbox{ if }\alpha<\Omega_{i+1}\\ 1&\mbox{ if }\alpha=\Omega_{i+1},\end{array}\right$.
  2. 2.
    $\chi^{\Omega_{i+1}}(\alpha):=\chi^{\Omega_{i+1}}(\eta)$ if $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$,
  3. 3.
    $\chi^{\Omega_{i+1}}(\alpha):=\left\{\begin{array}[]{cl}\chi^{\Omega_{i+1}}(\Delta)&\mbox{ if }\eta\not\in\mathrm{Lim}\mbox{ or }F_{j}(\Delta,\eta)\\[5.69054pt] \chi^{\Omega_{i+1}}(\eta)&\mbox{ otherwise,}\end{array}\right$.
    if $\alpha=\vartheta_{j}(\Delta+\eta)>\Omega_{i+1}$ and hence $j\geq i+1$.
For $\alpha\in{\operatorname{T}}$ we define $\mathrm{d}(\alpha):=i+1$ if $\chi^{\Omega_{i+1}}(\alpha)=1$ for some $i<\omega$ and $\mathrm{d}(\alpha):=0$ otherwise, so that
${\mathring{{\operatorname{T}}}}=\{\alpha\in{\operatorname{T}}\mid\mathrm{d}(\alpha)=0\}$
is the subset of $\operatorname{T}$} containing the terms of ordinals of (at most) countable cofinality.
Note that we obtain a partitioning of $\operatorname{T}$} in terms of cofinality through the preimages $\mathrm{d}^{-1}(n)$, $n<\omega$, cf. Lemma 3.4 of [3]. We separately define the support term $\underline{\alpha}$ of a $\vartheta$-term $\alpha$ first, as follows.
Definition 2.2 (cf. 3.5 of [3])
Let $\alpha\in{\operatorname{T}}$ be of the form $\alpha=\vartheta_{i}(\Delta+\eta)$ for some $i<\omega$, and denote the $\Omega_{i}$}-localization of $\alpha$ by ${\Omega_{i}}=\alpha_{0},\ldots,{\alpha_{m}}=\alpha$. The support term $\underline{\alpha}$ for $\alpha$ is defined by
[…]
Note that this includes setting the support term for $\vartheta_{0}(0)=1$ to 0, and, more importantly, that $\underline{\alpha}$ need not be a subterm of $\alpha$, namely when $\eta$ is a successor ordinal. However, we may consider $\underline{\alpha}$ to enter the inductively generated set of terms (e.g. $\operatorname{T}$}) before $\alpha$ since it is either 0, a subterm of $\alpha$, or it requires one generative step less (addition of $1$) when $\eta$ is a successor.
Definition 2.3 (cf. 3.5 of [3])
Let $\alpha\in{\operatorname{T}}$. By recursion on the build-up of $\alpha$ we define the function
$\alpha[\cdot]:\aleph_{d}\to{\operatorname{T}}^{\Omega_{d}}$
where $d:=\mathrm{d}(\alpha)$. Let $\zeta$ range over $\aleph_{d}$.
  1. 1.
    $0[\zeta]:=1[\zeta]:=0$.
  2. 2.
    $\alpha[\zeta]:=\xi+\eta[\zeta]$ if $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$.
  3. 3.
    For $\alpha=\vartheta_{i}(\Delta+\eta)>1$ where $i<\omega$, noting that $d\leq i$, the definition then proceeds as follows.
    1. 3.1.
      If $\eta\in\mathrm{Lim}$ and $\neg F_{i}(\Delta,\eta)$, that is, $\eta\in\mathrm{Lim}\cap\sup_{\sigma<\eta}\vartheta_{i}(\Delta+\sigma)$, we have $d=d(\eta)$ and define
      $\alpha[\zeta]:=\vartheta_{i}(\Delta+\eta[\zeta]).$
    2. 3.2.
      If otherwise $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$, we distinguish between the following 3 subcases.
      1. 3.2.1.
        If $\Delta=0$, define
        $\alpha[\zeta]:=\left\{\begin{array}[]{cl}\underline{\alpha}\cdot\zeta&\mbox{ if }\eta>0\mbox{ (and hence $d=0$)}\\ \zeta&\mbox{ otherwise.}\end{array}\right$.
      2. 3.2.2.
        $\chi^{\Omega_{i+1}}(\Delta)=1$. This implies that $d=0$, and we define recursively in $n<\omega$
        $\alpha[0]:=\vartheta_{i}(\Delta[\underline{\alpha}])\quad\mbox{ and }\quad\alpha[n+1]:=\vartheta_{i}(\Delta[\alpha[n]]).$
      3. 3.2.3.
        Otherwise. Then $d=d(\Delta)$ and
        $\alpha[\zeta]:=\vartheta_{i}(\Delta[\zeta]+\underline{\alpha}).$
Bachmann property for the above notation system holds in the restricted sense that for terms $\alpha,\beta$ such that $\alpha[\zeta]<\beta<\alpha$ and that are not regular cardinals, we have $\alpha[\zeta]\leq\beta[1]$, where $\zeta<\aleph_{d}$ and $d=\mathrm{d}(\alpha)$. Note first that the base cases of the recursive definition of fundamental sequences of limit ordinals are clauses 3.2.1, covering additive principal successors and regular cardinals, and 3.2.2, covering successor fixed points. The modification we made therefore only affects limits of countable cofinality the fundamental sequence of which is recursively based on clause 3.2.1. Clause 3.2.2 is the only case where a fundamental sequence is defined in terms of another fundamental sequence of different cofinality. Thus, given $\alpha$, $\beta$ such that $\alpha[\zeta]<\beta<\alpha$ and that are not regular cardinals, we still have $\alpha[\zeta]\leq\beta[0]$, unless the fundamental sequence for $\beta$ is recursively based on clause 3.2.1, in which case $\alpha[\zeta]\leq\beta[1]$ still holds.
2.2 Quotients
For any $\beta\in{\operatorname{T}}$ it can be syntactically detected whether $\beta$ is of a form $\beta=\alpha[n]$ for some $\alpha\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ and $n\in[2,\omega)$. Before we establish this explicitly, we use it in the following
Definition 2.4
We inductively define the elements of $({\nicefrac{{{\operatorname{T}}}}{{k}}})_{2\leq k<\omega}$ as increasing sequence of subsets of $\operatorname{T}$} as follows:
  1. 1.
    $0\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$.
  2. 2.
    If $\xi=_{\mathrm{\scriptscriptstyle{ANF}}}\xi_{1}+\ldots+\xi_{n}$ where $n\geq 0$ and $\eta\in{\mathbb{P}}$ such that $\eta<\xi_{n}$ if $n>0$, then $\xi,\eta\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$ implies that $\xi+\eta\cdot l\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$ if and only if $l<k$.
  3. 3.
    If $\xi\in{\nicefrac{{{\operatorname{T}}}}{{k}}}\cap\Omega_{i+2}$ where $i<\omega$, then $\beta:=\vartheta_{i}(\xi)\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$, unless:
    1. (a)
      $\underline{\beta}\not\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$, or
    2. (b)
      $\beta$ is of a form $\alpha[n]$ where $\alpha\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ and $n\geq k$.
The quotients ${\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}}\subseteq{\mathring{{\operatorname{T}}}}$ are then defined by ${\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}}:={\nicefrac{{{\operatorname{T}}}}{{k}}}\cap{\mathring{{\operatorname{T}}}}$, and we define the term sets relevant for the Goodstein process by
[…]
Remark 2.5
It is easy to see that we have
[…]
By means of the following inversion lemma we will be able to recover $\beta$ from terms $\alpha\in{\operatorname{T}}$ of a form $\beta[n]$ where $n\in[2,\omega)$ and $\beta\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, as used in Definition 2.4. As in Definition 3.5 of [3] we need to formulate the inversion lemma more generally, so as to cover terms of uncountable cofinality. The inversion is formulated in a way that will cover the cases needed to define the quotients of Definition 2.4, rather than covering all possible cases.
Note that we rely on the modified version Definition 3.5 of [3] of fundamental sequences for $\operatorname{T}$}, redefining $\eta\cdot\omega[n]:=\eta\cdot n$ instead of $\eta\cdot(n+1)$ for $\eta\in{\mathbb{P}}$ as mentioned in the introduction.
We first introduce a partial auxiliary function that will smoothen the formulation of the subsequent lemma, cf. the support term $\underline{\alpha}$ in Definition 2.2.
Definition 2.6
For $i<\omega$, $\Delta\in{\operatorname{T}}$ such that $\Omega_{i+1}\mid\Delta<\Omega_{i+2}$, and $\rho\in{\operatorname{T}}$ we define the partial function ${\eta}(i,\Delta,\rho)$ as follows.
[…]
Remark 2.7
In the cases where $\eta:={\eta}(i,\Delta,\rho)$ is defined, for $\alpha:=\vartheta_{i}(\Delta+\eta)$ we then have $\underline{\alpha}=\rho$, and either $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$ holds.
Lemma 2.8 (Inversion Lemma)
Let $\alpha\in{\operatorname{T}}$. $\alpha$ is of a form $\alpha=\beta[\zeta]$ where
  1. 1.
    $\zeta\in[2,\omega)$ and $\beta\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, or
  2. 2.
    $\zeta\in[{\Omega_{i}},\Omega_{i+1})\cap{\mathbb{P}}^{>1}$ for some $i<\omega$ and $\beta\in{\operatorname{T}}$ such that $\chi^{\Omega_{i+1}}(\beta)=1$
if and only if one of the following cases applies:
  1. 1.
    $\alpha=\zeta\in[{\Omega_{i}},\Omega_{i+1})\cap{\mathbb{P}}^{>1}$ and $\beta=\Omega_{i+1}=\vartheta_{i+1}(0)$.
  2. 2.
    $\alpha=\eta\cdot n$ for some $\eta\in{\operatorname{T}}\cap{\mathbb{P}}$ and $\zeta=n\in[2,\omega)$. Then $\beta:=\eta\cdot\omega$, and we have $\underline{\beta}=\eta$ and $\alpha=\beta[\zeta]$.
  3. 3.
    $\alpha=_{\mathrm{\scriptscriptstyle{ANF}}}\xi_{1}+\ldots+\xi_{k}+\rho$ for some $k\geq 1$ and $\rho=\eta[\zeta]$ according to the lemma’s conditions where $\eta\in(1,\xi_{k}]\cap{\mathbb{P}}$ so that $\rho<\xi_{k}$. Then setting $\beta:=\xi_{1}+\ldots+\xi_{k}+\eta$ we have $\alpha=\beta[\zeta]$.
  4. 4.
    $\alpha=\vartheta_{j}(\Gamma+\rho)$ and one of the following subcases applies:
    1. (a)
      $\rho$ is of a form $\rho=\eta[\zeta]$ according to the lemma’s conditions and $F_{j}(\Gamma,\eta)$ does not hold. Then setting $\Delta:=\Gamma$ and $\beta:=\vartheta_{j}(\Delta+\eta)$ we have $\alpha=\beta[\zeta]=\vartheta_{j}(\Delta+\eta[\zeta])$.
    2. (b)
      Otherwise, setting $\xi_{1}:=\Gamma+\rho$, check whether there is a (shortest) sequence $\xi_{1},\ldots,\xi_{m+1}$ (where $m\geq 1$) that determines a term $\Delta$ with $\chi^{\Omega_{j+1}}(\Delta)=1$ in the first step, such that
      1. i.
        $\xi_{k}$ is of a form $\Delta[\vartheta_{j}(\xi_{k+1})]$ according to the lemma’s conditions (here according to condition 2. with i=j) for $k=1,\dots,m$, and
      2. ii.
        $\xi_{m+1}$ is of a form $\Delta[\nu]$ where $\eta:=\eta(j,\Delta,\nu)$ is defined, so that $\beta:=\vartheta_{j}(\Delta+\eta)$, $\underline{\beta}=\nu$, and $\alpha=\beta[m]$.
      This case then applies if $\zeta=m\geq 2$.
    3. (c)
      $\Gamma$ is of a form $\Delta[\zeta]$ according to the lemma’s conditions, so that $\alpha=\vartheta_{j}(\Delta[\zeta]+\rho)$ where $\eta:=\eta(j,\Delta,\rho)$ is defined, so that $\beta=\vartheta_{j}(\Delta+\eta)$, $\underline{\beta}=\rho$, and $\alpha=\beta[\zeta]$.
Proof. Correctness follows by induction on $\alpha\in{\operatorname{T}}$: If one of cases $1$ - $4$ holds with $\beta$ and $\zeta$, then $\alpha=\beta[\zeta]$ matches either condition $1$ or $2$. The reverse direction, completeness, follows by induction on $\beta\in{\operatorname{T}}$, showing that for any $\beta\in{\operatorname{T}}$ and $\zeta$ according to either condition $1$ or $2$, the ordinal $\alpha:=\beta[\zeta]$ satisfies one of cases $1$ - $4$. $\Box$
Remark 2.9
Note that the outcome in case 4 (b) might be that $m=1$ and $\alpha=\beta[1]$, so that $\alpha$ fails to match the form detected for. Considering the example $\varepsilon_{0}$ as in the introduction, $\varepsilon_{0}[2]=\omega^{\omega}$ is seen as the first iterative repetition, despite $\varepsilon_{0}[1]=\omega=\vartheta_{0}(\vartheta_{0}(0))$.
We are now going to show that the <<-order type of the set ${{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}={\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}}\cap\Omega_{1}$ (where $k\in[2,\omega)$) is $\omega$ and that $\mathrm{G}_{k}$} enumerates this set, preserving its <<-order. To this end, we first define the analogue $\mathrm{imc}$} (for iterative maximal coefficient) of the measure $\operatorname{mc}$} that was used to define quotients $\nicefrac{{{\mathcal{E}}}}{{k}}$} of $\mathcal{E}$}. Recall Definition 2.4.
Definition 2.10
The iterative maximal coefficient $\mathrm{imc}$} of terms in $\operatorname{T}$} is defined as follows.
  1. 1.
    ${\mathrm{imc}}(0):=0$.
  2. 2.
    If $\xi=_{\mathrm{\scriptscriptstyle{ANF}}}\xi_{1}+\ldots+\xi_{n}\in{\operatorname{T}}$ where $n\geq 0$ and $\eta\in{\operatorname{T}}\cap{\mathbb{P}}$ such that $\eta<\xi_{n}$ if $n>0$, then for $l\in(0,\omega)$
    ${\mathrm{imc}}(\xi+\eta\cdot l):=\max\{{\mathrm{imc}}(\xi),{\mathrm{imc}}(\eta),l\}.$
  3. 3.
    If $\alpha\in{\operatorname{T}}$ is of a form $\vartheta_{i}(\xi)$ then
    ${\mathrm{imc}}(\alpha):=\max\{{\mathrm{imc}}(\xi),{\mathrm{imc}}(\underline{\alpha}),n\},$
    where either $n\geq 2$ is maximal such that $\alpha$ is of a form $\alpha=\beta[n]$ where $\beta\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, or otherwise $n=1$.
Note that we then have ${\mathrm{imc}}(\alpha)=0$ if and only if $\alpha=0$. $\mathrm{imc}$} is weakly monotone with respect to the subterm relation, hence weakly increasing along localization sequences. But it also needs to be monotone with respect to support terms (which as mentioned before need not be subterms). Consider the ordinals $\varepsilon_{0}\cdot\omega$ and $\omega^{\omega+1}$, where $\mathrm{G}_{2}$} yields the result $8$, but the former is in ${\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[2]$}}}$} while the latter is not.
We now obtain the following characterization of the quotients $\nicefrac{{{\operatorname{T}}}}{{k}}$} and $\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}$}.
Lemma 2.11
For $k\in[2,\omega)$ we have
[…]
Proof. This follows immediately by induction along the involved inductive definitions. $\Box$
Lemma 2.12
For $\alpha\in{\operatorname{T}}\cap\mathrm{Lim}$ and $\zeta\in{\operatorname{T}}\cap\aleph_{\mathrm{d}(\alpha)}$ we have
${\mathrm{imc}}(\alpha[\zeta])\leq\max\{{\mathrm{imc}}(\alpha),{\mathrm{imc}}(\zeta)\}.$
In particular, for $\alpha\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ we have ${\mathrm{imc}}(\alpha[n])\leq\max\{{\mathrm{imc}}(\alpha),n\}$.
Proof. The lemma is proved by induction along Definition 2.3. The i.h. immediately yields the claim for additively decomposable limit ordinals. Now suppose that $\alpha=\vartheta_{i}(\Delta+\eta)>1$.
  1. 1.
    If $\eta\in\mathrm{Lim}$ and $\neg F_{i}(\Delta,\eta)$, that is, $\eta\in\mathrm{Lim}\cap\sup_{\sigma<\eta}\vartheta_{i}(\Delta+\sigma)$, we have $d=d(\eta)$ and $\alpha[\zeta]=\vartheta_{i}(\Delta+\eta[\zeta])$, and the i.h. for $\eta$ applies. The term $\underline{\alpha[\zeta]}$ is either 0, a subterm of $\alpha[\zeta]$, or $\eta[\zeta]$ is a successor ordinal. In this latter case we must either have $\eta=\xi+\omega$ for some $\xi\in\mathrm{Lim}\cup\{0\}$ along with $\zeta\in(0,\omega)$, or $\eta=\xi+{\Omega_{j}}$ for some $\xi$ such that ${\Omega_{j}}\mid\xi$ and $j\leq i$ along with $\zeta$ a successor ordinal (represented in $\operatorname{T}$}), say $\zeta=\nu+l+1$ for some $\nu\in\mathrm{Lim}\cup\{0\}$ and $l<\omega$. Note that then $\alpha[\zeta]=\vartheta_{i}(\Delta+\xi+\nu+\omega)[l+1]$ and $l+1\leq{\mathrm{imc}}(\zeta)$. We obtain $\underline{\alpha[\zeta]}=\vartheta_{i}(\Delta+\xi+\zeta-1)$, so that ${\mathrm{imc}}(\underline{\alpha[\zeta]})\leq\max\{{\mathrm{imc}}(\alpha),{\mathrm{imc}}(\zeta)\}$.
  2. 2.
    If otherwise $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$, the definition of $\alpha[\zeta]$ distinguishes between the following 3 subcases.
    1. (a)
      If $\Delta=0$, we have
      $\alpha[\zeta]=\left\{\begin{array}[]{cl}\underline{\alpha}\cdot\zeta&\mbox{ if }\eta>0\mbox{ (and hence $d=0$)}\\ \zeta&\mbox{ otherwise,}\end{array}\right$.
      so that in case of $\eta=0$ the claim is obvious, and in case of $\eta>0$ we have $\alpha[\zeta]=\underline{\alpha}\cdot\zeta$ where $\zeta<\omega$.
    2. (b)
      $\chi^{\Omega_{i+1}}(\Delta)=1$. So we have $d=0$, $\zeta<\omega$, and according to the recursive definition
      $\alpha[0]:=\vartheta_{i}(\Delta[\underline{\alpha}])\quad\mbox{ and }\quad\alpha[n+1]:=\vartheta_{i}(\Delta[\alpha[n]])\mbox{ for }n<\omega.$
      We need to verify that ${\mathrm{imc}}(\underline{\alpha[\zeta]})\leq\max\{{\mathrm{imc}}(\alpha),{\mathrm{imc}}(\zeta)\}$, which immediately follows by side induction on $\zeta$ since in general $\Delta[\xi]$ cannot be a successor for $\xi$ either 0 or an additive principal number.
    3. (c)
      Otherwise. Then $d=d(\Delta)$ and
      $\alpha[\zeta]=\vartheta_{i}(\Delta[\zeta]+\underline{\alpha}).$
      Here $\Delta[\zeta]$ is a proper multiple of $\Omega_{i+1}$ since $\chi^{\Omega_{i+1}}(\Delta)=0$. Thus, the i.h. smoothly applies to $\Delta$ and we obtain the claim. $\Box$
Lemma 2.13
For $\lambda\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ and $\beta\in{\mathring{{\operatorname{T}}}}\cap[\lambda[k],\lambda)$ where $k\geq 2$ we have ${\mathrm{imc}}(\beta)\geq k$ and hence
${\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}}\cap[\lambda[k],\lambda)=\emptyset.$
Proof. We first observe that by Lemma 2.12, for any $\beta\in{\mathring{{\operatorname{T}}}}$
${\mathrm{imc}}(\beta)<k\Rightarrow{\mathrm{imc}}(\beta[i])<k\mbox{ for any }i<k.$
Suppose now that for some $\lambda\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, $\beta\in[\lambda[k],\lambda)$ is least such that ${\mathrm{imc}}(\beta)<k$. This implies that $\beta\in\mathrm{Lim}$ such that $\lambda[k]<\beta<\lambda$, as ${\mathrm{imc}}(\lambda[k])\geq k$ and clearly ${\mathrm{imc}}(\gamma)\leq{\mathrm{imc}}(\gamma+1)$ for any $\gamma$. Bachmann property yields $\lambda[k]\leq\beta[1]$, and since ${\mathrm{imc}}(\beta[1])<k$, we cannot have $\lambda[k]<\beta[1]$ because of the minimality of $\beta$. But we cannot have $\lambda[k]=\beta[1]$ either, as this would imply that ${\mathrm{imc}}(\beta[1])\geq k$. $\Box$
The following theorem is in correspondence with Lemma 1 and Remarks 1 and 2 of [1].
Theorem 2.14
For $k\in[2,\omega)$ the restriction of $\mathrm{G}_{k}$} to the set ${\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}$} is an order isomorphism with image $\mathbb{N}$}.
Proof. This now follows by <<-induction on $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$ via possibly iterated application of Lemma 2.13. $\Box$
3 Goodstein process
Definition 3.1
Let $N\in{\mathbb{N}}$ and $k\in[2,\omega)$. According to Theorem 2.14 there exists a unique $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$ such that $N={\mathrm{G}_{k}}(\alpha)$, the base-$k$ representation of $N$. For any $l\in(k,\omega)$ the change of base from $k$ to $l$ for $N$ is defined by
$N[k\mapsto l]:={\mathrm{G}_{l}}(\alpha).$
The change of base $k$ to $\omega$ for $N$ is defined by
$N[k\mapsto\omega]:=\alpha.$
Now that Theorem 2.14 is established, Cichon’s trick applies as layed out in the introduction, and we obtain the reformulation of the generalized Goodstein principle (3). Since the function defined in (4) is expressed in terms of the Hardy hierarchy as in (6), if the generalized Goodstein principle were provable in $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$, we would obtain a proof of the totality of $H_{\alpha}$ for $\alpha:={\operatorname{T}}\cap\Omega_{1}$ in $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$ using the canonical fundamental sequence $\alpha[n]:=\vartheta_{0}(\ldots\vartheta_{n}(0)\ldots)$, which is completely contained in each ${\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}$}, so that $H_{\alpha}(k)=H_{\alpha[k]}(k)$, despite the well-known fact that $H_{\alpha}$ is not provably total in $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$. We thus obtain
Corollary 3.2
The Goodstein process defined according to base transformation as in Definition 3.1 always terminates, which however exceeds the proof-strength of the theory $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$.
We close with a characterization of the ordinals collected in the sets ${\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}$} in terms of maximality.
Remark 3.3 (Maximality of elements of ${\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.79996pt}{$\scriptstyle[k]$}}}$})
Any ordinal $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$ is maximal among all $\beta\in{\operatorname{T}}\cap\Omega_{1}$ such that ${\mathrm{G}_{k}}(\alpha)={\mathrm{G}_{k}}(\beta)$. This is seen as follows: assuming otherwise let $\beta>\alpha$ be minimal such that ${\mathrm{G}_{k}}(\alpha)={\mathrm{G}_{k}}(\beta)$. Then $\beta[k]<\alpha<\beta$, so that according to Lemma 2.13 it would follow that ${\mathrm{imc}}(\alpha)\geq k$, which is not the case as $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$.
Acknowledgement
I would like to thank Professor Andreas Weiermann for numerous insightful and motivating discussions on ordinal notation systems and Goodstein principles, and for the suggestion to verify the characterization in terms of maximality in Remark 3.3.